### v1

@@ 6 Conclusion
The code for the MFEMiFSI plugin will be maintained in the publicly accessible FEBio Plugin Repository and associated Git repositories (links coming soon). The example problems considered in this work can be found in the FEBio Model Repository at https://repo.febio.org:443/permalink/project/135. All discussions in this work refer to the state of the plugin at the time of submission. We anticipate improvements to the framework after submission; however, readers can always restore the repository to this state using Git.
@@
Author contributions Ryan T. Black: Writing - original draft, Writing – review & editing, Methodology, Investigation, Formal analysis, Software, Visualization, Conceptualization. Steve A. Maas: Writing - original draft, Writing – review & editing, Methodology, Investigation, Formal analysis, Software, Visualization, Conceptualization. Wensi Wu: Methodology, Investigation, Formal Analysis, Conceptualization, Supervision. Jalaj Maheshwari: Investigation, Methodology, Writing – review & editing Tzanio Kolev: Writing – review & editing, Supervision, Conceptualization. Jeffrey A. Weiss: Writing – review & editing, Supervision, Software, Funding acquisition. Matthew A. Jolley: Writing – original draft, Writing – review & editing, Supervision, Resources, Project administration, Funding acquisition, Conceptualization.
@@
Data availability The code for the MFEMiFSI plugin will be maintained at the FEBio Plugin Repository (link coming soon) and associated Git repositories (link coming soon). Example problems from this work are available at the FEBio Model Repository (https://repo.febio.org:443/permalink/project/135). Simulation data will be made available upon request.
@@
Acknowledgements This work was funded by The Cora Topolewski Pediatric Valve Center at Children’s Hospital of Philadelphia, the Topolewski Endowed Chair in Pediatric Cardiology, a CHOP Cardiac Center Innovation Grant, NIH R01 HL153166(MAJ), R01 GM083925 (SAM, JAW), T32 HL007915 (RTB) and K25 HL168235 (WW). This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 (LLNL–JRNL–2014818).

### v2

@@ 6 Conclusion
The code for the MFEMiFSI plugin will be maintained in the publicly accessible FEBio Plugin Repository and associated Git repository https://github.com/JolleyLab/FEBioMFEMiFSI. The example problems considered in this work can be found in the FEBio Model Repository at https://repo.febio.org:443/permalink/project/135. All discussions in this work refer to the state of the plugin at the time of submission. We anticipate improvements to the framework after submission; however, readers can always restore the repository to this state using Git.
@@
Author contributions RTB and SAM contributed equally to this work. RTB: Writing - original draft, Writing – review & editing, Methodology, Investigation, Formal analysis, Software, Visualization, Conceptualization. SAM: Writing - original draft, Writing – review & editing, Methodology, Investigation, Formal analysis, Software, Visualization, Conceptualization. WW: Methodology, Investigation, Formal Analysis, Conceptualization, Supervision. JM: Investigation, Methodology, Writing – review & editing TK: Writing – review & editing, Supervision, Conceptualization. JAW: Writing – review & editing, Supervision, Software, Funding acquisition, Conceptualization. MAJ: Writing – original draft, Writing – review & editing, Supervision, Resources, Project administration, Funding acquisition, Conceptualization.
@@
Data availability The code for the MFEMiFSI plugin will be maintained at the FEBio Plugin Repository and associated Git repository https://github.com/JolleyLab/FEBioMFEMiFSI. Example problems from this work are available at the FEBio Model Repository https://repo.febio.org:443/permalink/project/135. Simulation data will be made available upon request.
@@
Acknowledgements This work was funded by The Cora Topolewski Pediatric Valve Center at Children’s Hospital of Philadelphia, the Topolewski Endowed Chair in Pediatric Cardiology, a CHOP Cardiac Center Innovation Grant, NIH R01 HL153166(MAJ), R01 GM083925 (SAM, JAW), T32 HL007915 (RTB) and K25 HL168235 (WW). This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 (LLNL–JRNL–2014818). High‑performance computing resources provided by the Anvil supercomputer at Purdue University through an NSF ACCESS Discover grant (MCH250067) and by the Children’s Hospital of Philadelphia are gratefully acknowledged.